On a given afternoon, the probability that a certain manager will be in her office is 0.48 and the probability that she will be in her home is 0.27. Assuming that the manager’s office and home are in two different locations, find the probability that she will be neither in her office nor in her home on a given afternoon. 



1
Expert's answer
2021-05-10T16:44:41-0400

Probability in office P(O)=0.48

Probability in home P(H)=0.27


Probability that she will be neither in her office nor in her home on a given afternoon-


"=1-P(O')P(H')\\=1-0.48\times 0.27\\=1-0.1296\\=0.8704"